### v2

# Modeling Diffusion and Permeation Across the Stratum Corneum Lipid Barrier

Abstract
Human skin oils are a major sink for ozone in densely occupied indoor environments. Understanding how the resulting volatile and semivolatile organic oxidation products influence indoor air chemistry requires accurate representations not only of their emission into indoor air but also of their transport across the outermost skin barrier, the stratum corneum. Using molecular dynamics simulations, we investigate the passive permeation of acetone, 6-methyl-5-hepten-2-one, and water—two representative products of skin-oil oxidation and a reference compound—through a model stratum corneum lipid membrane. We determine position-dependent diffusivities using two complementary analyses based on the same set of simulations and evaluate their accuracy through a propagator analysis. The two approaches provide upper and lower bounds for the true diffusivity, which, when combined with previously reported free-energy profiles, yield permeabilities relevant for modeling macroscopic skin transport. Our results show that permeation is governed primarily by energetic barriers rather than by molecular mobility, and that the predicted transport coefficients vary by about one order of magnitude depending on the chosen diffusivity estimator. These findings provide molecular-level constraints for parameters used in indoor air chemistry models and establish a transferable framework for linking atomistic transport mechanisms to large-scale simulations of human exposure and indoor air quality.
[…]
{tocentry}[Uncaptioned image]
1 Introduction
In 2024, the mean global near-surface temperature exceeded the pre-industrial average by more than $1.5\text{\,}\mathrm{K}$,WorldMeteorologicalOrganizationWMO.WorldMeteorologicalOrganizationWMO.2025.StateGlobalClimate underscoring the increasingly tangible impacts of global warming on human health, behavior, and daily life.Rocque.Witteman.2021.HealthEffectsClimate, IntergovernmentalPanelOnClimateChangeIPCC.IntergovernmentalPanelOnClimateChangeIPCC.2022.Impacts15degCGlobal, Abdul-Nabi.Zahran.2025.ClimateChangeIts A plausible behavioral response—particularly among vulnerable populations such as older adults and individuals with pre-existing health conditions—is to spend more time indoors to mitigate thermal discomfort during periods of extreme heat. However, empirical evidence remains limited regarding how climate change is reshaping patterns of time use.White-Newsome.ONeill.2012.ClimateChangeHealth, Connolly.Connolly.2018.ClimateChangeAllocation
Nevertheless, because humans spend most of their time indoors, the impacts of climate change are largely mediated through indoor environments.InstituteOfMedicine.InstituteOfMedicine.2011.ClimateChangeIndoor, Spengler.Spengler.2012.ClimateChangeIndoor Accordingly, understanding how rising outdoor temperatures influence indoor air quality has become an increasingly important research priority.Nazaroff.Nazaroff.2013.ExploringConsequencesClimate, Mansouri.Blondeau.2022.ImpactClimateChange Models and simulations are expected to play a central role in evaluating the long-term effects of climate change on indoor environments.Zhao.Schieweck.2024.LongtermPredictionEffects, Zhao.Hussein.2025.LongtermPredictionClimate However, these modeling efforts rely on accurate physicochemical representations of the complex surface–air chemistry occurring indoors.VonDomaros.Tobias.2025.MolecularDynamicsSimulations
In this study, we address this challenge by investigating the permeation of two representative volatile and semivolatile organic compounds (VOCs) through human skin. The compounds of interest—acetone and 6-MHO (Figure 1)—are formed during the ozonolysis of skin lipids, which act as the dominant ozone scavengers in densely occupied indoor environments and thus constitute an important source of secondary indoor pollutants.Wisthaler.Weschler.2010.ReactionsOzoneHuman, Weschler.Weschler.2016.RolesHumanOccupant Kinetic models have been developed to describe the mass transport and chemical transformations of these oxidation products across the skin, surrounding air, and clothing.Lakey.Shiraiwa.2017.ChemicalKineticsMultiphase, Lakey.Shiraiwa.2019.ImpactClothingOzone However, the reliability of such models depends critically on accurate physicochemical parameters describing the partitioning of semivolatile compounds—such as 6-MHO—between the human body and indoor air,VonDomaros.Tobias.2020.MultiscaleModelingHuman as well as on reliable descriptions of their diffusional transport through the skin barrier.
To address this challenge, we recently began investigating permeation energetics using atomistic molecular dynamics (MD) simulations of stratum corneum (SC) lipid membranes (Fig 1)—the skin’s primary barrier to chemical penetration.thomas_insights_2025 We found that the convergence of free-energy (FE) profiles describing the translocation of permeants across these membranes was remarkably slow, requiring microseconds of simulation time even when state-of-the-art enhanced sampling techniques were employed. We traced the molecular origin of this convergence bottleneck to the highly ordered structure of SC membranes, which contrasts sharply with the fluid-like organization of conventional lipid bilayers such as those composed of 1-palmitoyl-2-oleoyl-sn-glycero-3-phosphocholine (POPC). In addition, we identified stochastic lipid flip–flop events between the two membrane leaflets that induce long-lived structural asymmetries.
Figure 1: Chemical structures of acetone and 6-MHO, and a representative snapshot of the model system. The atomistic lipid membrane is composed of cholesterol (orange), lignoceric acid (yellow), and ceramide NS (green). In this study, we investigate solute permeation across the membrane along the $z$-axis. Figure adapted with permission from ref. thomas_insights_2025.
In this work, we extend our previous investigations by providing estimates of diffusivity profiles along the membrane normal. Predicting these diffusivities remains a challenging task—particularly when permeation is slow and enhanced sampling is required to overcome large FE barriers. We employ two complementary approaches to compute position-dependent diffusivities from harmonically restrained MD simulations, based on either the velocity autocorrelation function (VACF)Woolf.Roux.1994.ConformationalFlexibilityOPhosphorylcholine or the position autocorrelation function (PACF)Hummer.Hummer.2005.PositiondependentDiffusionCoefficients of the coordinate describing membrane permeation. As shown by our results—and consistent with the findings of Rowley et al. for a dipalmitoylphosphatidylcholine (DPPC) membranegaalswyk_generalized_2016—estimates obtained from the two approaches differ substantially. Through a rigorous propagator-based analysis, we demonstrate that the resulting diffusivity profiles provide lower and upper bounds to the true profile. We further propagate these results to obtain single-membrane permeabilities and we critically assess how methodological uncertainties affect the accuracy of these predictions. Taken together, our work advances the understanding of molecular permeation across skin lipid membranes and establishes a robust framework for improving predictive models of indoor air chemistry under changing environmental conditions.
2 Methods
We use atomistic MD simulations, which follow the motion of all atoms in the system in time according to classical mechanics. This approach allows us to quantify how individual molecules move through a model skin–lipid barrier and how readily they cross it. The resulting microscopic transport properties serve as input parameters for larger-scale models of skin transport and indoor air chemistry. The model systems and simulation protocol have been described in detail in our previous study.thomas_insights_2025 Here, we briefly summarize the main aspects relevant to the present work and describe the computation of position-dependent diffusivities and the propagator analysis.
2.1 Model Systems
The SC is a complex structure composed of cornified cells embedded in a lipid matrix, an arrangement often described as a brick-and-mortar architecture.Michaels.Shaw.1975.DrugPermeationHuman, Elias.Elias.1983.EpidermalLipidsBarrier, Schmitt.Neubert.2020.StateArtStratum Atomistic studies of skin permeation typically focus on the lipid matrix, which is the only continuous component of the SC and represents a major barrier to permeation. Extrapolation to the full SC requires combining MD-based transport properties with macroscopic models that account for additional SC components and the detailed permeation pathway.Wang.Nitsche.2007.MultiphaseMicroscopicDiffusion, Wang.Nagel.2023.TransdermalLateralEffective
The structure of the SC lipid matrix has been extensively characterized using a range of experimental techniques. There is broad consensus that it consists predominantly of an approximately equimolar mixture of ceramides (CERs), free fatty acids (FFAs), and cholesterol (CHOL), organized into at least two lamellar microphases: the short periodicity phase (SPP) and the long periodicity phase (LPP).Sparr.Topgaard.2023.StratumCorneumBarrier, Bouwstra.Gooris.2023.SkinBarrierExtraordinary
[…]
Here, we investigate the permeation of three solutes—acetone, 6-MHO, and water—representing two skin-oil oxidation products and a reference molecule. These solutes permeate through a model SC lipid bilayer originally proposed by Wang and Klaudawang_models_2018 and representative of the SPP. The present work focuses on the SPP as a tractable and widely used model system. Establishing robust methodology for diffusivity estimation in this simpler context provides a foundation for future studies that will incorporate the LPP and more realistic SC architectures.
2.2 Umbrella Sampling
The translocation of the solutes across the membrane is described by a collective variable (CV), denoted as $z$, which measures the center-of-mass distance between each solute and the membrane projected onto the membrane normal (Figure 1). The CV $z$ can be viewed as a one-dimensional reaction coordinate that tracks how far the molecule has moved from the membrane center toward either side. Instead of relying on a single long simulation to capture rare permeation events, we divide the coordinate into many overlapping windows along $z$, where each window focuses the sampling on a narrow region of the membrane, in a procedure known as umbrella sampling (US). The solute is confined within each window by a harmonic potential, which permits fluctuations around a chosen position while still allowing the solute to explore the local environment. For each solute, we performed 91 US windows, in which the solute was harmonically restrained at integer-$\mathrm{\text{Å}}$ positions along the $z$ coordinate.
A detailed analysis of the influence of the harmonic force constant was presented by Rowley et al.,gaalswyk_generalized_2016 and our own related findings on this topic are provided in Section S1 of the Supporting Information (SI). For the results reported in the main text, a force constant of $30\text{\,}\mathrm{\text{kcal}}\text{\,}{\mathrm{mol}}^{-1}\text{\,}{\mathrm{\text{Å}}}^{-2}$ is employed. Each simulation is initiated from well-equilibrated configurations obtained in our previous study and run for $15\text{\,}\mathrm{ns}$ per window. Although simulation times per window are short, the combined sampling over all windows effectively corresponds to many independent permeation attempts across the entire membrane. All simulations are carried out at a temperature close to that of the outer skin surface ($T=$305.15\text{\,}\mathrm{K}$\$) to mimic realistic physiological conditions. The temperature is maintained by a stochastic velocity-rescaling thermostatbussi_canonical_2007 with a time constant of $1\text{\,}\mathrm{ps}$. All simulations are performed using NAMD 3.0 (alpha13).phillips_scalable_2005, phillips_scalable_2020
2.3 Diffusivity Profiles
In addition to the FE barriers, a key quantity for transport is how quickly a molecule can move at different depths within the membrane. We express this by a position-dependent diffusivity $D(z)$, which may be interpreted as a local mobility: large $D(z)$ indicates that the molecule can move easily at that position, whereas small $D(z)$ reflects strong hindrance by the surrounding lipids. Two approaches are used to compute the position-dependent diffusivities, $D(z)$, along the membrane normal.
The first approach, proposed by Woolf and Roux, relies on the calculation of the velocity autocorrelation function (VACF):Woolf.Roux.1994.ConformationalFlexibilityOPhosphorylcholine
$C_{v}(t)=\left<\dot{z}(t)\dot{z}(0)\right$>. (1)
The VACF quantifies how quickly the solute’s velocity loses memory of its initial direction of motion. If a molecule continues to move in approximately the same direction for a long time, it is effectively more mobile than if its velocity becomes randomized on a short timescale. To extract a position-dependent diffusivity $D(z)$, one must perform a Laplace transform of the VACF, $\hat{C}_{v}(s)$, and extrapolate the resulting Laplace-frequency–dependent diffusivity,
$D(s)=\frac{-\hat{C}_{v}(s)\left<\delta z^{2}\right>\left<\dot{z}^{2}\right>}{\hat{C}_{v}(s)\!\left[s\!\left<\delta z^{2}\right>+s^{-1}\!\left<\dot{z}^{2}\right>\right]-\left<\delta z^{2}\right>\left<\dot{z}^{2}\right>},$ (2)
to the limit of zero frequency (infinite lag time),
$D(z=\left<z\right>)=\lim_{s\to 0}D(s).$ (3)
However, this limit is ill-defined because of the $s^{-1}$ dependence and therefore requires an empirical extrapolation procedure.
Hummer later proposed an alternative approach that avoids Laplace-domain extrapolation by using the position autocorrelation function (PACF):Hummer.Hummer.2005.PositiondependentDiffusionCoefficients
$C_{\delta z}(t)$ $=\frac{\left<\delta z(t)\delta z(0)\right>}{\left<\delta z^{2}\right>},$ (4)
[…]
The PACF characterizes how strongly the current position of the solute along $z$ remains correlated with its initial position. In regions where the molecule moves freely, this positional correlation decays rapidly, whereas in regions where its motion is constrained, the correlation persists for longer times. In practice, however, one must either impose a cutoff time for the integral or assume a functional form to fit the correlation decay—both of which can introduce systematic bias.
The velocities required for computing the VACF are obtained from finite differences of the CV trajectory, which is saved at $2\text{\,}\mathrm{fs}$ intervals. Position-dependent diffusivities are determined using the ACFCalculator code developed by Rowley et al., which implements a heuristic procedure for the extrapolation in eq. 3.gaalswyk_generalized_2016
We note that alternative strategies for estimating position-dependent diffusivities are available. In particular, Bayesian-inference methods, which were originally proposed by Hummerhummer_position-dependent_2005 and later extended to anisotropic media by Ghysels et al.Ghysels.Hummer.2017.PositionDependentDiffusionTensors, have also been developed. They have also been coupled to adaptive biasing force calculations by Comer, Chipot, and co-workers.Comer.Gonzalez-Nilo.2013.CalculatingPositionDependentDiffusivity, Chipot.Comer.2016.SubdiffusionMembranePermeation Complementary formulations have been developed by Rosta and collaborators.Rosta.Hummer.2015.FreeEnergiesDynamic, Sicard.Rosta.2021.PositionDependentDiffusionBiased Position-dependent diffusivities can also be estimated from the autocorrelation function of the constraint force (FACF) in simulations where the particle is algorithmically constrained to selected positions $z$ along the CV. However, this approach can be sensitive to finite-size and momentum-conservation artifacts and therefore requires particular care in its practical implementation.FujimotoNagaiEtAl2021 We focus here on the VACF/PACF approaches because they are widely used in membrane permeation studies,das_water_2009, gupta_molecular_2016, kramer_membrane_2020, piasentin_evaluation_2021, Reuter.Lunter.2025.PresenceDifferentCeramide and they allow direct comparison with prior results for our benchmark system (water).
Using both VACF- and PACF-based estimators on the same simulation data allows us to assess how sensitive the inferred diffusivities are to the choice of analysis method. As shown below, the two approaches agree well in the bulk water regions but diverge near the membrane center, where the dynamics is more complex. We interpret these differences as providing a realistic range (lower and upper bounds) for the true local diffusivity.
2.4 Propagator Analysis
Diffusion in inhomogeneous media can be described by the Smoluchowski equation. In the one-dimensional case relevant to this study, the position- and time-dependent probability density $p(z,t)$ evolves according to
$\partial_{t}p(z,t)=\partial_{z}\left\{D(z)e^{-\beta\Delta F(z)}\,\partial_{z}\!\left[e^{\beta\Delta F(z)}p(z,t)\right]\right\},$ (6)
where $\beta=1/(k_{\mathrm{B}}T)$ is the inverse thermal energy, $\Delta F(z)$ denotes the FE profile, and $D(z)$ represents the position-dependent diffusivity. This one-dimensional Smoluchowski equation describes diffusion on a free-energy landscape: the molecule tends to move from regions of high free energy to regions of lower free energy, while its motion is locally slowed down or accelerated by the position-dependent diffusivity $D(z)$. In this formulation, the thermodynamic driving forces are encoded in $\Delta F(z)$ and the kinetic hindrance in $D(z)$, and both contributions enter a single evolution equation for the probability of finding the molecule at a given position and time.
For given $\Delta F(z)$ and $D(z)$ profiles and a delta-function initial condition, $p(z,t_{0})=\delta(z-z_{0})$, this equation can be solved numerically using a standard forward-in-time, centered-in-space (FTCS) finite-difference scheme. The resulting solution, $p(z,t\,|\,z_{0},t_{0})$, represents the propagator of the diffusion model—that is, the conditional probability of finding a particle initially located at $z(t_{0})=z_{0}$ at position $z(t)$ at a later time $t$. The same quantity can also be obtained directly from molecular dynamics (MD) simulations without assuming diffusive behavior, thereby providing a rigorous test of the accuracy of the proposed $\Delta F(z)$ and $D(z)$ profiles. Comparing these model-based propagators to those obtained directly from MD simulations provides a stringent test of the validity of the chosen $\Delta F(z)$ and $D(z)$ profiles.
We compute MD-derived propagators for the critical membrane-core region as follows. A total of 1800 snapshots were selected in which the solute was located near the membrane center ($|z|<$0.05\text{\,}\mathrm{\text{Å}}$\$). The harmonic restraint was then removed, allowing the solute to freely explore the entire membrane. Unrestrained simulations were run for $15\text{\,}\mathrm{ns}$, and propagators $p(z,t\,|\,z_{0}=0,t_{0}=0)$ for selected lag times $t$ were estimated from histograms along the $z$ coordinate.
3 Results and Discussion
3.1 Diffusivity Profiles
Figure 2: Transmembrane diffusivity profiles obtained from the velocity autocorrelation function (VACF) and position autocorrelation function (PACF) analyses. Quantitative and qualitative discrepancies are evident near the membrane center. Solid lines are shown to guide the eye, and dashed lines indicate reference values for the aqueous-phase bulk diffusivities. In the water profile, two representative data points—one in the bulk region (black) and one at the membrane center (red)—are highlighted for further analysis.
Diffusivity profiles, $D(z)$, of the solutes along the transmembrane axis were estimated from the same set of simulations using the PACF and VACF approaches described above. The resulting profiles are shown in Figure 2.
[…]
In the headgroup region, both methods indicate a local diffusivity bottleneck for all solutes. The most pronounced discrepancy between the two approaches, however, appears at the membrane center, where the estimates differ both quantitatively and qualitatively: the VACF method predicts higher-than-bulk diffusivities, whereas the PACF method yields lower values. Similar observations were reported by Rowley et al. for a DPPC membrane.gaalswyk_generalized_2016 In the following, we discuss the challenges associated with determining diffusivities using both approaches and the likely origins of these discrepancies. Our analysis focuses primarily on two representative regions of the membrane: the aqueous region, where the two methods show good agreement, and the membrane center, where the discrepancies are most pronounced.
3.2 Challenge 1: Extrapolation of the VACF Results
Figure 3: Illustration of determining diffusivities from velocity autocorrelation functions. The quantity $D(s)$, computed from eq. 2, must be extrapolated to zero by fitting a straight line over a range of $s$ values where the function remains well-behaved. This procedure performs reliably in the bulk phase but is highly sensitive to the placement of the tangent line for windows at the membrane center. The values obtained using the automated extrapolation procedure of Rowley et al.gaalswyk_generalized_2016 are indicated by the black and red points in Figure 2.
Obtaining position-dependent diffusivities from VACFs requires computing a Laplace-frequency–dependent diffusivity, $D(s)$, for each US window (eq. 2), followed by extrapolation to zero frequency (eq. 3). However, this limit is ill-defined because of the $s^{-1}$ dependence of $D(s)$ and therefore necessitates an empirical extrapolation procedure. Our VACF-based diffusivity profiles were obtained using the automated linear extrapolation method of Rowley et al.gaalswyk_generalized_2016 Although this approach is relatively robust in the aqueous bulk region, it may not be optimal at the membrane center (Figure 3). Because $D(s)$ exhibits pronounced curvature in the membrane core, the linear extrapolation is highly sensitive to the placement of the tangent line, yielding diffusivity estimates between roughly $0.1\text{\,}\mathrm{\text{Å}}^{2}\mathrm{/}\mathrm{p}\mathrm{s}$ and $0.75\text{\,}\mathrm{\text{Å}}^{2}\mathrm{/}\mathrm{p}\mathrm{s}$.
3.3 Challenge 2: Determination of PACF Correlation Times
Figure 4: Illustration of determining diffusivities from position autocorrelation functions (PACFs). Shown is the estimated diffusivity, $D(t)$, as a function of the PACF integration cutoff time in two selected regions of the membrane. In the absence of noise, $D(t)$ should reach a plateau. A clear plateau is visible in the bulk water region but not at the membrane center. The values corresponding to a fixed cutoff of $t=$10\text{\,}\mathrm{ps}$\$ are indicated by the black and red points in Figure 2.
The challenges associated with extrapolating $D(s)$ to zero frequency were recognized by Hummer, who proposed an alternative approach based on determining the PACF correlation time (eq. 5).hummer_position-dependent_2005 For purely diffusive dynamics, this method is formally equivalent to the VACF approach, but it is considerably simpler to apply in practice.
In Figure 4, we show selected diffusivities, $D(t)$, as a function of the integration cutoff time used in eq. 5. The PACF method performs well when the solute resides in the aqueous phase, where $D(t)$ reaches a clear plateau between approximately $8\text{\,}\mathrm{ps}$ and $10\text{\,}\mathrm{ps}$. Although noise leads to some fluctuations, it does not significantly alter the predicted diffusivity. In contrast, at the membrane center, the decay of the PACF is substantially slower, and no clear plateau in $D(t)$ is observed even after $10\text{\,}\mathrm{ps}$. These observations are consistent with those reported by Rowley et al. for a DPPC membrane.gaalswyk_generalized_2016
As a consequence, solute- and region-specific integration cutoffs would be ideal but require tedious manual adjustment. Instead, the diffusivity profiles in Figure 2 were obtained using practical compromise cutoffs: $8\text{\,}\mathrm{ps}$ for water and $10\text{\,}\mathrm{ps}$ for acetone and 6-MHO. Although these fixed cutoffs may slightly overestimate PACF-derived diffusivities in certain regions (e.g., the membrane center), our main conclusions are unaffected: PACF-derived profiles still represent lower bounds to the true $D(z)$.
3.4 Propagator Analysis
Figure 5: Free-energy profiles for the translocation of all solutes across the model stratum corneum membrane, obtained from umbrella sampling (US) and well-tempered metadynamics (WTM).thomas_insights_2025 The two methods produce consistent overall features, with only minor quantitative differences. Figure adapted with permission from ref. thomas_insights_2025.
Figure 6: Propagators, $p(z,t\,|\,z_{0},t_{0})$, representing the probability of finding a particle initially located at $z_{0}=0$ at position $z$ after a lag time $t$. Shown are results obtained by numerically solving the Smoluchowski diffusion equation using diffusivity profiles derived from the velocity autocorrelation function (VACF) and position autocorrelation function (PACF) analyses, compared with direct simulation data. At short lag times ($t=$2\text{\,}\mathrm{ps}$\$), the VACF-based model shows quantitative agreement with the simulation data, although the true diffusive regime has likely not yet been reached on this timescale. At longer lag times ($t=$100\text{\,}\mathrm{ps}$\$), the VACF- and PACF-based models provide upper and lower bounds, respectively, to the true propagators and, by extension, to the diffusivity profile.
Having obtained two qualitatively distinct diffusivity profiles from the same set of simulations, it is natural to ask which of the two, if either, is correct. To address this question, we computed the propagators of both diffusion models in the critical membrane-center region, $p(z,t\,|\,z_{0}=0,t_{0}=0)$—that is, the probability of finding a particle initially located at $z_{0}=0$ at position $z$ after a lag time $t$. The corresponding FE profiles were taken from our previous study,thomas_insights_2025 as shown in Figure 5. For the present analysis, we used the umbrella-sampling (US)–based free-energy profiles. A comparison between the model-derived propagators and those obtained directly from molecular dynamics simulations is presented in Figure 6 for two representative lag times.
At short lag times ($t=$2\text{\,}\mathrm{ps}$\$), the VACF-based model shows quantitative agreement with the simulation data over six orders of magnitude, whereas the PACF approach fails to do so. This finding is consistent with the results reported by Rowley et al.,gaalswyk_generalized_2016 who performed a simplified analysis by approximating the propagator as a normal distribution—an assumption that holds only for lag times long enough to reproduce diffusive dynamics yet short enough that the effects of inhomogeneity remain negligible.
However, it is not obvious that the lag time of $2\text{\,}\mathrm{ps}$ chosen here (or the $5\text{\,}\mathrm{ps}$ used by Rowley et al.) is sufficiently long for the diffusive approximation to hold. Indeed, the failure of the PACF to decay within such short timescales may also be interpreted as a breakdown of this assumption. Subdiffusive behavior has, in fact, been reported for lag times up to at least $64\text{\,}\mathrm{ps}$ for methanol permeation across a POPC membrane.Chipot.Comer.2016.SubdiffusionMembranePermeation
We therefore extended our propagator analysis to selected lag times of up to $500\text{\,}\mathrm{ps}$. For a range of lag times between $50\text{\,}\mathrm{ps}$ and $500\text{\,}\mathrm{ps}$, which we consider to lie within the truly diffusive regime, the model propagators exhibit the behavior shown for the representative case of $t=$100\text{\,}\mathrm{ps}$\$ in Figure 6. The complete set of results is provided in Section S3.
At these timescales, neither the VACF- nor the PACF-based models reproduce the simulation data quantitatively. However, the VACF results consistently overestimate the diffusive spread (i.e., the diffusivities are too large), whereas the PACF results underestimate it (i.e., the diffusivities are too small). We therefore interpret the two profiles as providing upper and lower bounds, respectively, to the true diffusivity profile, which must lie between them. In the following, these bounds are used to assess the sensitivity of derived quantities, specifically the transmembrane permeabilities.
3.5 Impact on Permeabilities
Figure 7: Membrane permeabilities estimated using diffusivity profiles derived from the velocity autocorrelation function (VACF, left) and position autocorrelation function (PACF, right) analyses, each combined with free-energy profiles obtained from umbrella sampling (US) and well-tempered metadynamics (WTM). Differences between the free-energy methods are noticeable only for acetone, whereas discrepancies between the two diffusivity analyses consistently span approximately one order of magnitude across all solutes.
Within the inhomogeneous solubility–diffusion (ISD) model,Diamond.Katz.1974.InterpretationNonelectrolytePartition, Marrink.Berendsen.1994.SimulationWaterTransport the local resistance of a membrane to permeation, $R(z)$, is expressed in terms of the diffusivity and FE profiles as
$R(z)=\frac{e^{\beta\Delta F(z)}}{D(z)}\,.$ (7)
The overall permeability of the membrane, $P$, is then obtained by integrating the resistance over the membrane width, $h$, according to
$\frac{1}{P}=\int_{-h/2}^{+h/2}R(z)\,dz.$ (8)
Using the FE profiles for the translocation of all solutes across the membrane (Figure 5), we computed the membrane permeabilities according to eqs 7 and 8. The resulting permeabilities are shown in Figure 7, and the corresponding numerical values are provided in the Supporting Information (Section S4). Because of the exponential dependence on the free energy in eq 7, even small variations in $\Delta F(z)$ are amplified, leading to pronounced differences in the resulting permeabilities, which follow the order 6-MHO >> acetone $\gg$ water.
We have also shown that the FE profiles obtained from US and WTM differ slightly (Figure 5), even after extensive sampling, although these differences remain within the limits of statistical uncertainty.thomas_insights_2025 With the diffusivity profiles now available, we can propagate the different FE profiles to evaluate their impact on the resulting transport coefficients. We find good agreement between the results obtained from both methods for water and 6-MHO; for acetone, however, noticeable deviations are observed. Nevertheless, the permeability coefficients for all solutes differ by less than an order of magnitude, indicating that statistical variations between FE methods likely have little to no impact on kinetic model predictions.
More striking are the differences in permeabilities when comparing VACF- and PACF-derived diffusivities. Because the VACF-derived diffusivities are consistently higher than those obtained from the PACF method, the same trend is reflected in the corresponding permeabilities. In absolute terms, deviations reach approximately one order of magnitude. Whether such differences are significant for kinetic models is highly problem specific. In many cases, these discrepancies are likely acceptable, although the final assessment should be guided by a dedicated sensitivity analysis.Lakey.Shiraiwa.2019.ImpactClothingOzone
It is important to note here that diffusivity, free energy, and permeation reflect different aspects of transport. A high free-energy barrier (as in the case of water at the bilayer center) reduces the probability of occupancy in that region but does not necessarily imply fast diffusion once a molecule is there. Permeability depends on both thermodynamics and kinetics through eq. 7, and intuition based solely on an unfavourable free-energy environment risks conflating the role of $\Delta F(z)$ with that of $D(z)$. Diffusivity profiles $D(z)$ probe the mobility of a solute conditioned on being at position $z$, not the likelihood of being there in the first place. Within this interpretation, a lower diffusivity $D(z)$ at the membrane center does not mean that water prefers this region, only that its motion is more hindered when present. A surprising consequence of this interplay between thermodynamic and kinetic effects is the relatively high permeability of 6-MHO, despite the fact that its diffusivity profile is generally the lowest of the three solutes.
[…]
Figure 8: Local resistance, $R(z)$, estimated using free-energy profiles obtained from umbrella sampling and diffusivity profiles derived from the velocity autocorrelation function analysis. The corresponding results for other combinations of free-energy and diffusivity methods are provided in Section S4. The red, green, and gray vertical lines indicate different plausible integration limits. The areas under the curves in these regions are negligible compared with the contributions from the membrane center, indicating that different integration limits have little impact on the overall permeability.
To investigate the impact of these definitions, we show the local resistance $R(z)$ in Figure 8. The area under each curve corresponds to the inverse permeability, $P^{-1}$. Within the membrane interior, this area is several orders of magnitude larger than the small additional contributions from the interfacial regions. We therefore conclude that different reasonable integration limits have little to no effect on the computed permeabilities.
We conclude this section by noting that a direct molecular interpretation of the resistance profiles in Fig. 8 in terms of the membrane structure in Fig. 1 is non-trivial. The local resistance $R(z)$ reflects the combined effect of free-energy barriers and diffusivity (eq. 7), meaning that structural features do not map one-to-one onto the profile but instead contribute through a convolution of thermodynamic and kinetic influences. Nevertheless, we can conjecture about some qualitative structure–transport relations. The elevated resistance near the headgroup region is consistent with the dense packing and strong hydrogen-bonding environment at lipid–water interfaces, which restricts mobility particularly for polar permeants. In the membrane center, the loosened lipid organization and resulting elevated free volume give rise to conditions that may accelerate local diffusion and contribute to the central resistance dip.
Table 1: Water permeability coefficients of stratum corneum (SC) lipid membranes, obtained from molecular dynamics simulations and experimental compilations. Simulation results are reported together with the corresponding force fields, temperatures, and analysis methods used to compute permeability, which differ across studies and contribute to variability in the reported values. Units are $\mathrm{cm}\text{\,}{\mathrm{s}}^{-1}$.
Reference Permeability Approach Details
Gupta et al. gupta_molecular_2016
$7.08\pm 1.18\times 10^{-5}$ FACF 310 K, GROMOS/SPC
Das et al. das_water_2009
$8.2\times 10^{-8}$ FACF 350 K, OPLS/SPC
Reuter et al. reuter_presence_2025
$1.88\times 10^{-5}$ FACF 303.15 K, CHARMM36/TIP3P
Piasentin et al. piasentin_evaluation_2021
$2.79\pm 0.52\times 10^{-5}$ FACF 303.15 K, CHARMM36/TIP3P
$2.10\pm 0.39\times 10^{-5}$ US + FACF 303.15 K, CHARMM36/TIP3P
$1.89\pm 0.34\times 10^{-5}$ US + VACF 303.15 K, CHARMM36/TIP3P
This work
$5.65\times 10^{-5}$ US + PACF 305.15 K, CHARMM36/TIP3P
$5.48\times 10^{-5}$ WTM + PACF 305.15 K, CHARMM36/TIP3P
$4.15\times 10^{-4}$ US + VACF 305.15 K, CHARMM36/TIP3P
$4.00\times 10^{-4}$ WTM + VACF 305.15 K, CHARMM36/TIP3P
Katritzky et al. katritzky_skin_2006a
$2.95\times 10^{-7}$ Experimental
Chen et al. chen_modeling_2010a
$4.79\times 10^{-7}$ Experimental
  • a
    The reported permeability values are compiled from multiple secondary sources and do not represent primary experimental measurements.
4 Conclusions
We used MD simulations to quantify position-dependent diffusivities, $D(z)$, for water, acetone, and 6-MHO across a representative SC lipid membrane and combined these results with previously reported FE profiles to compute transmembrane permeabilities, $P$. By applying both PACF- and VACF-based analyses to the same set of US trajectories and validating them through a propagator comparison, we demonstrated that the two $D(z)$ estimates systematically bracket the true diffusivity profile in the diffusive regime, with PACF providing a lower bound and VACF an upper bound. Permeabilities derived from these bounds consequently span approximately one order of magnitude, whereas differences arising from the choice of FE method (US vs. WTM) become negligible once propagated to transport parameters. Consistent with the exponential dependence of the ISD framework on free energy, the ordering of permeabilities follows the underlying FE landscape rather than the local diffusivities—6-MHO >> acetone $\gg$ water—highlighting that permeation across the membrane is governed primarily by energetic barriers rather than by differences in molecular mobility.
To place these results in context, we compare the water permeability coefficients obtained in this work with values reported in previous studies. The permeabilities derived from PACF-based diffusivities are of the same order of magnitude as those reported by Gupta et al., Piasentin et al., and Reuter et al. gupta_molecular_2016, piasentin_evaluation_2021, Reuter.Lunter.2025.PresenceDifferentCeramide, namely on the order of $10^{-5}$ $\mathrm{cm}\text{\,}{\mathrm{s}}^{-1}$ (see Table 1). Residual discrepancies are most likely attributable to differences in simulation temperature and force-field parametrization. Additional methodological factors—such as system size, thermostat and barostat settings, integration parameters, and overall simulation protocols—may also contribute to variations in the reported permeability coefficients. Importantly, this level of agreement should not be interpreted as a validation of the PACF-based diffusivity profiles. The limitations of the approach discussed here are methodological in nature and are therefore expected to be largely independent of specific computational details.
Direct comparison of permeabilities computed from our MD simulations to experimental values (in vitroanderson_heterogeneity_1988, anderson_solute_1989 or in vivokalia_homogeneous_1996) is not straightforward because our model represents a simplified SC membrane containing only the short periodicity phase (SPP) of the lipid matrix. We deliberately omitted the more structurally complex long periodicity phase (LPP) to first address methodological challenges that arise even in comparatively simple SC models. Improving the realism of atomistic representations is important, but equally essential is ensuring that these methodological limitations are understood and managed if future permeability estimates are to be more robust. We therefore emphasize the need to quantify uncertainty—whether through error bars, bounding estimates, or credibility intervals—and to evaluate how these propagate into higher-scale models.
Permeability coefficients are often estimated empirically using quantitative structure–activity relationship (QSAR) and continuum diffusion models,wilschut_estimating_1995, lian_evaluation_2008, Mitragotri.Roberts.2011, Frasch.Barbero.2013, Wang.Naegel.2023 which correlate measured permeabilities with molecular descriptors such as molar mass, partition coefficients, and bulk diffusion constants. Although such models are powerful for interpolation within their training domains, they do not explicitly capture the molecular-scale interactions and dynamic heterogeneity that govern transport across the SC lipid matrix—especially when applied to molecules that fall outside their original parameterization range. In contrast, MD simulations capture these effects directly by resolving how solutes interact with and move through the heterogeneous lipid environment. For example, despite its larger molar mass relative to water and acetone, 6-MHO exhibits the highest permeability because of its comparatively low free-energy barriers to permeation (Figure 5), underscoring the dominant role of thermodynamic driving forces—rather than molecular size—in determining SC transport.
With respect to indoor air chemistry, several kinetic models—such as KM-BL, KM-FILM, and KM-SUB-Skin-ClothingLakey.Shiraiwa.2019.ImpactClothingOzone—have been developed to describe the dynamics of skin-oil oxidation products across different interfaces. However, these models critically depend on physicochemical parameters that are difficult or impossible to measure experimentally, underscoring the value of molecular-scale simulations like those presented here. For example, Lakey et al. demonstrated that the performance of their KM-FILM model is highly sensitive to the assumed diffusion constants of surface films,lakey_kinetic_2021 illustrating how improved microscopic input can directly enhance macroscopic predictive accuracy.
Taken together, our results provide mechanistic and quantitative constraints—via upper and lower bounds on $D(z)$ and permeabilities—that can be directly used to refine existing indoor-air kinetic models. Beyond these applications, the present framework establishes a foundation for systematically linking atomistic transport processes to macroscopic exposure and air-quality models. Future directions include (i) extending the approach to multilamellar SC architectures and anisotropic membrane transport; (ii) assessing the influence of force-field and water-model choice; (iii) developing more robust and transferable diffusivity estimators; and (iv) performing targeted experimental validation for VOCs of indoor relevance. Collectively, these efforts will help bridge the scale gap between molecular transport and human exposure in complex indoor environments.
{acknowledgement}
The authors thank Chinmay Das and Peter D. Olmsted for helpful discussions on setting up SC models. DJT acknowledges support from the Alfred P. Sloan Foundation Chemistry of the Indoor Environment (CIE) Program (G-2020-13912).
{suppinfo}
[…]